OpenStudy (anonymous):

A baseball player hits a ball. The ball leaves the bat with an initial upward velocity of 35 meters per second. If the ball is 1 meter off the ground when it leaves the bat, about how many seconds will it take for the ball to hit the ground?

OpenStudy (anonymous):

i hate these

OpenStudy (saifoo.khan):

speed = distance over time.

OpenStudy (anonymous):

thats nt the answer

OpenStudy (anonymous):

ah now we have a problem!

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

was it hit vertically up in the air?

OpenStudy (anonymous):

do you know what formula to use?

OpenStudy (anonymous):

7 seconds -0.03 seconds 14.34 seconds 35.133 seconds

OpenStudy (anonymous):

its one of those

OpenStudy (saifoo.khan):

0.03 (not in nedative)

OpenStudy (saifoo.khan):

(negative)

OpenStudy (anonymous):

uhh negatie is the only one

OpenStudy (anonymous):

you have to use the formula for free fall

OpenStudy (anonymous):

wtf idk how thats why im asking u guys lol

myininaya (myininaya):

do you have an example for another problem with a solution?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i know it in feet. myininaya knows it in meters

OpenStudy (anonymous):

i think it is \[-4.9t^2\] yes?

OpenStudy (anonymous):

in meters that is. so you have to solve \[1+35t-4.9t^2=0\] for t. lets see mr khan do that in two seconds

OpenStudy (anonymous):

in any case i think you get 7

OpenStudy (anonymous):

finally i got one!

OpenStudy (anonymous):

.3s

OpenStudy (anonymous):

Vy= dy/ t . solve for t.. t= dy/Vy, amnd thus 1/35, seconds cancel out, .3m

OpenStudy (anonymous):

wouldn't you need to know the initial upward acceleration of the ball at the moment it left the bat?

OpenStudy (anonymous):

let me get my physics referance table, is 1m its max height

OpenStudy (anonymous):

oh, you are right. ignore my posts.

OpenStudy (anonymous):

these problems are nonsense don't take them seriously

OpenStudy (anonymous):

they are pretending that the ball goes straight up in the air. the whole exercise is pointless

OpenStudy (anonymous):

and so to all a \[\huge{\color{red}{goodnight!}}\]